# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ 12.3. Integer crossing obstruction
Case 4: $s=3$, $d\geq 3$. The all-inert bound gives $G(3)\leq\zeta(3)^{2}/\zeta(6)=\prod_{p}(p^{3}+1)/(p^{3}-1)$. The partial product through $p=5$ equals $(9\cdot 28\cdot 126)/(7\cdot 26\cdot 124)=567/403$. Each remaining factor satisfies $(p^{3}+1)/(p^{3}-1)<1+3p^{-3}$ (since $2/(p^{3}-1)<3/p^{3}$ for $p\geq 2$), so $\prod_{p>5}(p^{3}+1)/(p^{3}-1)<\exp(3\sum_{p>5}p^{-3})<\exp(3/50)$, using $\sum_{p>5}p^{-3}<\int_{5}^{\infty}x^{-3}\,dx=1/50$. Since $\exp(y)<1+y+y^{2}$ for $0<y<1$, the tail is bounded by $\exp(3/50)<2659/2500$, giving $G(3)\leq(567/403)(2659/2500)=1507653/1007500$. Then $1507653^{2}=2{,}273{,}017{,}568{,}409<3{,}045{,}168{,}750{,}000=3\cdot 1007500^{2}$, so $G(3)<\sqrt{3}\leq\sqrt{d}$.
@@
$G(3)^{2}\leq\frac{178326^{2}}{138229^{2}}=\frac{31{,}800{,}162{,}276}{19{,}107{,}256{,}441}<2,$
since $31{,}800{,}162{,}276<2\times 19{,}107{,}256{,}441=38{,}214{,}512{,}882$. ∎
@@ 13.4. Weight robustness
$N_{w}$ 10.02-10.02 8.630-8.630 21.08-21.08 99.37-99.37
Heat kernel $e^{-0.1\gamma^{2}}$ $S_{\zeta}(w)$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$ $3.5\times 10^{-9}$
@@ A.1. Software and rigorous arithmetic
Computations in this paper fall into two categories. Rigorous computations use interval arithmetic via ARB [1], accessed through python-flint $\geq 0.8.0$ (Python 3.12); ARB represents numbers as balls $m\pm r]$ with midpoint $m$ and radius $r$, and arithmetic operations propagate error bounds rigorously, so the true value is guaranteed to lie within the output ball as a mathematical theorem. This includes all zero certification, spectral sum bounds, Besicovitch constants, and Bessel product integrals (via acb_calc_integrate). The PSLQ integer relation search uses mpmath $\geq 1.3$; the Ferguson–Bailey–Arno bounds [3] ensure that None results rigorously certify the absence of integer relations with bounded coefficients (see §A.7).
Working precisions: 512-bit ($\approx 154$ decimal digits) for $S_{\zeta}^{*}$ certification, spacelike verification (Table 2), and null crossing computation (ARB); 500-bit for the auxiliary $L$-function zeros tabulated in Appendices C and E (20 decimal places, computed with an earlier version of the zero-finding pipeline); 1500-bit ($\approx 451$ decimal digits) for all $\chi_{5}$ zero refinement and certification (ARB); 128-bit for the sign-change scan and bracket filtering phases described in §A.2, used in the extended 1000-zero computations of Appendix F; 256-bit for all Bessel product integrals including density, resonance, and decay estimates (ARB); and 80-digit decimal precision for PSLQ searches (mpmath).
@@ A.2. $L$-function evaluation and zero-finding
$\theta_{\chi}(t)=\begin{cases}\operatorname{Im}\!\bigl(\log\Gamma(s/2)\bigr)-\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=+1\text{ (even)},\\[4.0pt] \operatorname{Im}\!\bigl(\log\Gamma((s{+}1)/2)\bigr)-\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=-1\text{ (odd)},\end{cases}$
@@
Phases 1 and 2: Interleaved sign-change scan and bracket filtering. Evaluate $Z_{\chi}(t)$ on a uniform grid with step $0.05$ over $0,T_{\max}]$, where $T_{\max}$ is chosen so that the zero-counting formula $N(T,\chi_{d})\sim\frac{T}{\pi}\log\frac{qT}{2\pi e}$ exceeds the number of zeros sought by a safety margin; for example, $200$ zeros of $L(s,\chi_{5})$ require $T_{\max}\approx 290$. All arithmetic in this phase uses 128-bit ARB precision. At each sign change $Z_{\chi}(t_{a})\,Z_{\chi}(t_{b})<0$, $20$ bisection steps are performed immediately (interleaved with the grid scan), reducing the bracket width from $0.05$ to $0.05\cdot 2^{-20}\approx 5\times 10^{-8}$ and producing a seed accurate to approximately six decimal digits. The bisection simultaneously filters spurious sign changes arising from Hardy-phase oscillations: at the bisected midpoint, $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ is evaluated and the bracket is accepted as a genuine zero only if $|L|<10^{-4}$; brackets with $|L|>10^{-3}$ are discarded. This interleaved design avoids storing all sign changes before filtering, so scanning stops as soon as the required number of genuine seeds has been collected.
Phase 3: Newton refinement. Refine each seed by Newton iteration applied to $Z_{\chi}(t)$. The derivative is approximated by the finite difference $(Z_{\chi}(t+h)-Z_{\chi}(t))/h$; the midpoint of the resulting ARB ball is extracted after each Newton step (see Remark A.1 below). Two modes are used depending on the target precision:
  • Standard mode (128-bit, $h=10^{-8}$): from a six-digit seed, quadratic convergence reaches $\sim$20 correct digits in two to three iterations. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-18}$.
  • High-precision mode (1500-bit, $h=10^{-50}$): from a six-digit seed, seven to eight iterations reach full working precision ($\approx 451$ digits), with certified bounds on $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ typically in the range $10^{-358}$$10^{-450}$. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-68}$.
@@
Certification. A zero ordinate $\gamma^{\prime}$ is accepted when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<\text{threshold}$, with the threshold depending on the target precision ($10^{-18}$ for 20-digit values, $10^{-68}$ for 70-digit values). The certified bound is computed by evaluating $L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})$ in ARB ball arithmetic and extracting the upper endpoint of the absolute value ball; this is a rigorous upper bound as a mathematical theorem.
Precomputation. The constants $\log q$, $\log(q/\pi)$, and the Hurwitz parameters $a/q$ for each nonzero character value are computed once and stored as arb objects at the working precision. This avoids redundant string-to-ball conversions in the inner loop and reduces the per-evaluation cost by approximately $30\%$.
@@
Three points are essential for correct implementation.
(1) Newton iteration must target $Z_{\chi}$, not $L$. On the critical line, $L(\tfrac{1}{2}+it,\chi_{d})$ is complex-valued for non-trivial characters; its real part does not vanish at zeros of $L$, so Newton iteration on $\operatorname{Re}L$ converges to a spurious fixed point rather than to a zero. The Hardy phase rotation reduces the problem to root-finding for a real function of one real variable, where Newton’s method achieves quadratic convergence (doubling the number of correct digits per iteration).
(2) Midpoint extraction is required in ball arithmetic. The Newton update $t\mapsto t-Z(t)/Z^{\prime}(t)$ introduces radius from the division. Without extracting the ball midpoint after each step, the radius grows geometrically and the computation produces NaN within a few iterations. This is specific to rigorous interval arithmetic and does not arise in ordinary floating-point computation.
(3) The finite-difference step $h$ must be appropriate for the current precision. The step $h=10^{-50}$ is suitable for 1500-bit ($\approx 451$-digit) working precision: it is large enough to avoid catastrophic cancellation (the difference Z(t+h)Z(t)Z(t+h)-Z(t) retains $\approx 400$ significant digits) and small enough to make the $O(h)$ truncation error negligible relative to the quadratic convergence. For 128-bit standard mode, $h=10^{-8}$ provides the analogous balance at lower precision. A step that is too small for the working precision (for example, $h=10^{-200}$ at 512-bit) produces a derivative estimate dominated by rounding error, causing Newton iteration to stall.
@@ A.5. Computation of $\epsilon_{M}$
The tail bounds $\epsilon_{M}=\sum_{k>M}b_{k}$ reported in the stability table of Section 7 were computed using the methodology of §A.1A.2. The first 200 zeros of $L(s,\chi_{5})$ were located and certified as described therein; the Lorentzian weights $b_{k}=2/({\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2}})$ were computed in ARB and summed. The analytic tail beyond zero 200 (at height $T=283.93$) was bounded by the integral $\frac{1}{\pi}\bigl(\frac{\log(5T/2\pi)}{T}+\frac{1}{T}\bigr)=0.0072$, derived from Abel summation against the zero-counting formula $N(T,\chi_{5})\sim\frac{T}{\pi}\log\frac{5T}{2\pi e}$.
@@ A.6. Bessel product integrals
The density at the origin $f_{S_{L}}(0)$ and the mean absolute value $\mathbb{E}[|S_{\zeta}|]$ appearing in Corollary 9.5 are computed via rigorous numerical quadrature using ARB’s acb_calc_integrate function, which implements the Petras algorithm with adaptive Gauss-Legendre quadrature and certified error bounds [1]. All results are rigorous ARB balls.
@@
For the mean absolute value (16): $\mathbb{E}[|S_{\zeta}|]=\frac{2}{\pi}\int_{\epsilon}^{T}\frac{1-\prod_{k=1}^{20}J_{0}(a_{k}t)}{t^{2}}\,dt$ with $\epsilon=10^{-3}$ and T=50, [^0] T=50{,}000, where $a_{k}=2/(\tfrac{1}{4}+\gamma_{k}^{2})$ for $\zeta$ zeros. The integrand is $O(1)$ near $t=0$ (since $1-\prod J_{0}(a_{k}t)=O(t^{2})$); the contribution from $0,\epsilon]$ is bounded by $(\sum_{k}a_{k}^{2}/4)\cdot\epsilon<4\times 10^{-8}$ and added to the main integral. ARB quadrature yields:
$\mathbb{E}[|S_{\zeta}|]=0.00716132817045\ldots\;\pm\;2.20\times 10^{-24}.$
The product $C(5)=2\cdot f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ is certified to four significant figures.

### 12 Mar 2026 v3

@@ 77 $28$ $2/7$ $17$ $17$ $8/17$
Case 4: $s=3$, $d\geq 3$. The all-inert bound gives $G(3)\leq\zeta(3)^{2}/\zeta(6)=\prod_{p}(p^{3}+1)/(p^{3}-1)$. The partial product through $p=5$ equals $(9\cdot 28\cdot 126)/(7\cdot 26\cdot 124)=567/403$. Each remaining factor satisfies $(p^{3}+1)/(p^{3}-1)<1+3p^{-3}$ (since $2/(p^{3}-1)<3/p^{3}$ for $p\geq 2$), so $\prod_{p>5}(p^{3}+1)/(p^{3}-1)<\exp(3\sum_{p>5}p^{-3})<\exp(3/50)$, using $\sum_{p>5}p^{-3}<\int_{5}^{\infty}x^{-3}\,dx=1/50$. Since $\exp(y)<1+y+y^{2}$ for $0<y<1$, the tail is bounded by $\exp(3/50)<2659/2500$, giving $G(3)\leq(567/403)(2659/2500)=1507653/1007500\approx 1.496$. Then $1507653^{2}=2{,}273{,}017{,}568{,}409<3{,}045{,}168{,}750{,}000=3\cdot 1007500^{2}$, so $G(3)^{2}\approx 2.239<3$ and $G(3)<\sqrt{3}\leq\sqrt{d}$. An independent ARB certificate for this case, using the direct bound $(\zeta(3)^{2}/\zeta(6))^{2}<3$ without the partial-product decomposition, is provided in the ancillary scripts.
@@
$G(3)^{2}\leq\frac{178326^{2}}{138229^{2}}=\frac{31{,}800{,}162{,}276}{19{,}107{,}256{,}441}\approx 1.664<2,$
since $31{,}800{,}162{,}276<2\times 19{,}107{,}256{,}441=38{,}214{,}512{,}882$. An independent ARB certificate, computing $G(3,2)=\zeta(3)/L(3,\chi_{8})$ directly via Hurwitz decomposition and certifying $G(3,2)^{2}<2$, is provided in the ancillary scripts. ∎
@@ 13.4. Weight robustness
Remark 13.10 (Scope of the robustness result).
Proposition 13.9 establishes only the instantaneous spacelike property $N_{w}<0$. Three further statements in this paper are not claimed to hold for general admissible weights, and do require the specific Lorentzian structure:
  1. (i)
    Cesàro spacelike ($\langle N_{w}\rangle_{T}<0$ for all $T>0$, Theorem 4.10). This requires a three-region lower bound on the bilinear form $h(T)=\langle S_{L}^{2}\rangle_{T}$, which uses the exact values $b_{k}=2/({\tfrac{1}{4}}+\gamma_{k}^{\prime 2})$ and the specific interference structure of those weights. For a general admissible weight, analogous bounds would require detailed knowledge of the weight’s interaction with the zero-spacing statistics.
  2. (ii)
    Effective density bound ($\operatorname{dens}\{N_{d}>0\}\leq A_{M}/\sqrt{d}+B_{M}$, Theorem 8.5). This uses the finiteness of $\|f_{S_{L}^{(M)}}\|_{\infty}$, which is established via Bessel product analysis with the specific weights $b_{k}$. A general admissible weight produces a Cesàro distribution whose density at the origin may be infinite.
  3. (iii)
    Density rate and exact asymptotic ($O(1/\sqrt{d})$, Corollary 8.8; $C(d)/\sqrt{d}$, Theorem 9.4). These depend on the Jessen–Wintner structure of the Lorentzian Cesàro distribution and the specific computable constant $C(d)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]$.
The introduction states that “the spacelike property is robust to the choice of weight function”; this refers to conclusion (i) of the present proof, not to conclusions (ii) or (iii).
@@
$N_{w}$ 10.02-10.02 8.632-8.632 21.08-21.08 99.37-99.37
Heat kernel $e^{-0.1\gamma^{2}}$ $S_{\zeta}(w)$ $2.1\times 10^{-9}$ $2.1\times 10^{-9}$ $2.1\times 10^{-9}$ $2.1\times 10^{-9}$
@@ A.1. Software and rigorous arithmetic
Computations in this paper fall into two categories. Rigorous computations use interval arithmetic via ARB [1], accessed through python-flint $\geq 0.8.0$ (Python 3.12); ARB represents numbers as balls $m\pm r]$ with midpoint $m$ and radius $r$, and arithmetic operations propagate error bounds rigorously, so the true value is guaranteed to lie within the output ball as a mathematical theorem. This includes all zero certification, spectral sum bounds, Besicovitch constants, and Bessel product integrals (via acb.integral). The PSLQ integer relation search uses mpmath $\geq 1.3$; the Ferguson–Bailey–Arno bounds [3] ensure that None results rigorously certify the absence of integer relations with bounded coefficients (see §A.7).
Working precisions: 512-bit ($\approx 154$ decimal digits) for $S_{\zeta}^{*}$ certification, spacelike verification (Table 2), and null crossing computation (ARB); at least 1500-bit ($\approx 451$ decimal digits) for all zero refinement and certification (ARB), including both the $\chi_{5}$ zeros of Appendix B and the auxiliary character zeros of Appendix C; 128-bit for the sign-change scan and bracket filtering phases described in §A.2, used in the extended 1000-zero computations of Appendix F; 256-bit for all Bessel product integrals including density, resonance, and decay estimates (ARB); and 80-digit decimal precision for PSLQ searches (mpmath).
@@ A.2. $L$-function evaluation and zero-finding
$\theta_{\chi}(t)=\begin{cases}\operatorname{Im}\!\bigl(\log\Gamma(s/2)\bigr)+\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=+1\text{ (even)},\\[4.0pt] \operatorname{Im}\!\bigl(\log\Gamma((s{+}1)/2)\bigr)+\tfrac{t}{2}\log(q/\pi)&\text{if }\chi_{d}(-1)=-1\text{ (odd)},\end{cases}$
@@
Phases 1 and 2: Interleaved sign-change scan and bracket filtering. Evaluate $Z_{\chi}(t)$ on a uniform grid with step $0.05$ over $0,T_{\max}]$, where $T_{\max}$ is chosen so that the zero-counting formula $N(T,\chi_{d})\sim\frac{T}{\pi}\log\frac{qT}{2\pi e}$ exceeds the number of zeros sought by a safety margin of $2.2\times$, with the additional floor $T_{\max}\geq 2N$; for example, $200$ zeros of $L(s,\chi_{5})$ require $T_{\max}=400$. All arithmetic in this phase uses 128-bit ARB precision. At each sign change $Z_{\chi}(t_{a})\,Z_{\chi}(t_{b})<0$, $20$ bisection steps are performed immediately (interleaved with the grid scan), reducing the bracket width from $0.05$ to $0.05\cdot 2^{-20}\approx 5\times 10^{-8}$ and producing a seed accurate to approximately six decimal digits. The bisection simultaneously filters spurious sign changes arising from Hardy-phase oscillations: at the bisected midpoint, $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ is evaluated and the bracket is accepted as a genuine zero only if $|L|<10^{-4}$; all other brackets are discarded as spurious. This interleaved design avoids storing all sign changes before filtering, so scanning stops as soon as the required number of genuine seeds has been collected.
Phase 3: Newton refinement. Refine each seed by Newton iteration applied to $L(s,\chi_{d})$ on the critical line. Each Newton step computes $s_{n+1}=s_{n}-L(s_{n})/L^{\prime}(s_{n})$ and projects the result back to $\sigma=\tfrac{1}{2}$ by extracting the midpoint of the imaginary part. Both $L(s,\chi_{d})$ and $L^{\prime}(s,\chi_{d})$ are evaluated as certified ARB balls via acb_series.zeta, which computes the Taylor series of the Hurwitz zeta function $\zeta(s+x,a/q)$ to order $2$ at $x=0$; the two coefficients are $\zeta(s,a/q)$ and $\zeta^{\prime}(s,a/q)$ respectively. No finite-difference approximation is used. Two modes are used depending on the target precision:
  • Standard mode (128-bit): from a six-digit seed, quadratic convergence reaches $\sim$20 correct digits in two to three iterations. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-18}$.
  • High-precision mode (at least 1500-bit): from a six-digit seed, seven to eight iterations reach full working precision ($\approx 451$ digits), with certified bounds on $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ typically in the range $10^{-430}$$10^{-452}$. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-68}$. Zeros with bounds weaker than $10^{-400}$ are boosted to $2200$-bit (Phase 8).
@@
Certification. Each zero is certified by two independent mechanisms. First, the ARB ball $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<\text{threshold}$ provides a rigorous magnitude bound. Second, the winding number of $L$ around the thin rectangle $0.49,0.51]\times[\gamma^{\prime}-h,\gamma^{\prime}+h]$ (with $h=10^{-8}$, falling back to larger $h$ if the narrow box is undecidable at the working precision) is computed via the argument principle in ARB arithmetic. A winding number of $1$ certifies that exactly one zero of $L$ lies in the rectangle. Since $\chi_{d}$ is a real primitive character, the functional equation forces off-line zeros to occur in symmetric pairs $\rho,1-\bar{\rho}$; a rectangle of width $0.02$ centered on $\sigma=\tfrac{1}{2}$ cannot contain both members of such a pair unless both lie on the critical line, so the unique zero must have $\sigma=\tfrac{1}{2}$.
Phase 4: Global completeness verification (the seal). The argument principle counts zeros of $L(s,\chi_{d})$ in the wide rectangle $0.5,1.5]\times[t_{\mathrm{lo}},T_{\max}]$ via ARB-rigorous argument tracking around the four sides of the contour. If the winding number equals the number of certified zeros, the table is complete. As a secondary validation, the winding number count is compared against the analytic lower bound from the BMOR explicit formula (Bennett–Martin–O’Bryant–Rechnitzer [2]). This contour-integral approach to completeness verification follows Johansson (arXiv:1802.07942), who computed $N(T)$ for $\zeta(s)$ via a rectangular contour integral and noted that the result is “a ball that provably determines $N(T)$ as an integer.”
Phases 5 and 6: Recovery. If Phase 4 detects a discrepancy (more zeros by winding number than in the table), strip-by-strip winding checks identify which inter-zero gaps are deficient. For each deficient gap, a narrow rectangular winding number localizes the missing zero and its midpoint provides a seed. These recovery seeds are refined by Newton iteration at $1500$-bit (Phase 6), merged into the table, and Phase 4 is rerun on the repaired table. The final seal is the global contour count on the repaired table.
Precomputation. The constants $\log(q/\pi)$ and the Hurwitz parameters $a/q$ for each nonzero character value are computed once and stored as arb objects at the working precision. This avoids redundant string-to-ball conversions in the inner loop and reduces the per-evaluation cost by approximately $30\%$.
@@
Two points are essential for correct implementation.
(1) Newton iteration targets $L$, not $Z_{\chi}$. The Newton update $s_{n+1}=s_{n}-L(s_{n})/L^{\prime}(s_{n})$ uses the complex-valued $L$-function directly, with the real part projected back to $\sigma=\tfrac{1}{2}$ after each step. This is valid because the projection discards the (small) off-line component of the Newton step while retaining the imaginary-part correction. Working with $L$ and its exact derivative (via acb_series.zeta) avoids the Hardy phase rotation entirely during refinement; the Hardy $Z$-function is used only in Phases 1 and 2 for the sign-change scan.
(2) Midpoint extraction is required in ball arithmetic. The Newton update $s\mapsto s-L/L^{\prime}$ introduces radius from the division. Without extracting the ball midpoint after each step, the radius grows geometrically and the computation produces NaN within a few iterations. This is specific to rigorous interval arithmetic and does not arise in ordinary floating-point computation.
@@ A.5. Computation of $\epsilon_{M}$
The tail bounds $\epsilon_{M}=\sum_{k>M}b_{k}$ reported in the stability table of Section 7 were computed using the methodology of §A.1–§A.2. The first 200 zeros of $L(s,\chi_{5})$ were located and certified as described therein; the Lorentzian weights $b_{k}=2/({\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2}})$ were computed in ARB and summed. The analytic tail beyond zero 200 (at height $T=283.935$) was bounded by the integral $\frac{1}{\pi}\bigl(\frac{\log(5T/2\pi)}{T}+\frac{1}{T}\bigr)=0.0072$, derived from Abel summation against the zero-counting formula $N(T,\chi_{5})\sim\frac{T}{\pi}\log\frac{5T}{2\pi e}$.
@@ A.6. Bessel product integrals
The density at the origin $f_{S_{L}}(0)$ and the mean absolute value $\mathbb{E}[|S_{\zeta}|]$ appearing in Corollary 9.5 are computed via rigorous numerical quadrature using ARB’s acb.integral function, which implements the Petras algorithm with adaptive Gauss-Legendre quadrature and certified error bounds [1]. All results are rigorous ARB balls.
@@
For the mean absolute value (16): $\mathbb{E}[|S_{\zeta}|]=\frac{2}{\pi}\int_{\epsilon}^{T}\frac{1-\prod_{k=1}^{20}J_{0}(a_{k}t)}{t^{2}}\,dt$ with $\epsilon=10^{-30}$ and T=50, [^0] T=50{,}000, where $a_{k}=2/(\tfrac{1}{4}+\gamma_{k}^{2})$ for $\zeta$ zeros, plus the analytic correction $(2/\pi)/T$ for $\int_{T}^{\infty}t^{-2}\,dt$ and the negligible Bessel envelope tail (bounded below $10^{-17}$ in ARB). The integrand is $O(1)$ near $t=0$ (since $1-\prod J_{0}(a_{k}t)=O(t^{2})$); the Petras algorithm handles the removable singularity by adaptive subdivision. ARB quadrature at 256-bit precision yields:
$\mathbb{E}[|S_{\zeta}|]=0.00717406056590\ldots\;\pm\;2.58\times 10^{-62}.$
The product $C(5)=2\cdot f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1193$ is certified to four significant figures.
Unsigned Bessel product integrals. Several tables in Sections 67 report integrals of the form $\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{M}|J_{n_{k}}(b_{k}t)|\,dt$, where the absolute values introduce cusps at every zero of each Bessel factor. Because acb.integral (the Petras algorithm) assumes analyticity in a strip around the integration path, it is not directly applicable to these integrands. All unsigned integrals in this paper are therefore evaluated by strip decomposition: the domain $[0,T]$ is partitioned at the certified zeros of every Bessel factor into thin strips (of half-width $\Delta=10^{-20}$ in the argument $b_{k}t$, centered on each zero) and the gaps between them. On each gap the integrand is analytic and sign-definite, so acb.integral is rigorous. On each strip the integrand is bounded by the Landau constant $b=0.675$ (Landau [8], Theorem 1), giving a per-strip error of $O(\Delta)$ which is negligible ($10^{-17}$) in all cases. Bessel zeros are computed entirely in ARB via McMahon asymptotic seeds (DLMF 10.21.19) followed by certified trichotomy bisection; IVT, non-overlap, and certified ordering checks are performed for every zero. This architecture is shared by the ancillary scripts for the tables after Lemma 6.6, Remark 6.12, Lemma 7.1, Remark 7.2, and Lemma 7.3.